# HiPreNets: High-Precision Neural Networks through Progressive Training

### v2

@@ Abstract
Deep neural networks are powerful tools for solving nonlinear problems in science and engineering, but training highly accurate models becomes challenging as problem complexity increases. Non-convex optimization and numerous hyperparameters to tune make performance improvement difficult, and traditional approaches often prioritize minimizing mean squared error (MSE) while overlooking $L^{\infty}$ error, which is the critical focus in many applications. To address these challenges, we present a progressive framework for training and tuning high-precision neural networks (HiPreNets). Our approach refines a previously explored staged training technique for neural networks that improves an existing fully connected neural network by sequentially learning its prediction residuals using additional networks, leading to improved overall accuracy. We discuss how to take advantage of the structure of the residuals to guide the choice of loss function, number of parameters to use, and ways to introduce adaptive data sampling techniques. We validate our framework’s effectiveness through several benchmark problems.
@@ 1 Introduction
Fully-connected deep neural networks [26, 27, 15] have found widespread use in applications such as computer vision [18] and natural language processing [31, 24], yet they remain underutilized in scientific and engineering domains due to concerns about accuracy and safety [14]. Guaranteeing high accuracy on these problems is difficult because they are often high-dimensional and nonlinear. Much of the existing literature focuses on attaining a “good enough” average error rather than devoting significant effort to drive it closer to machine precision. Additionally, the final maximum error of a neural network approximation is generally ignored despite worst-case behavior being critical to analyze for sensitive applications [2].
@@
In recent work, [4, 20] attempted to extend boosting to neural networks by proposing the idea of boosted neural networks. In these methods, fully connected neural networks are used as the “weak” models to be combined. These models are less interpretable and more costly to optimize than decision trees, but they are much better at capturing nonlinear patterns, can produce robust outputs, and can more easily be regularized to avoid overfitting. In [33], this approach was taken one step further by recognizing neural networks’ tendency to learn the low-frequency components of functions first, known as the spectral bias of neural networks [25]. As such, the low and high frequency parts of a dataset and the models’ residuals can be analyzed and incorporated into the boosting algorithm to further guide the training process. This analysis is insightful and was shown to work for low dimensional problems, but it appears much more difficult to apply it to higher dimensional problems. In summary, while these earlier works have applied boosting concepts to neural networks, they typically focused on shallow models or spectral insights that are hard to generalize in practice, and none of them analyze methods for reducing worst-case error.
@@
To address these issues, we introduce a progressive training framework based on boosting in which the neural network is composed of multiple components trained sequentially. Each component has significantly lower complexity than the full network. This modularity of the framework also offers great flexibility by allowing for each component network to be trained with a different architecture and loss function that may better fit the data it encounters. By using a loss function and other training techniques at each iteration that incorporate information from the residuals, the training can be guided to preferentially reduce either the MSE or the maximum error.
@@
We focus exclusively on applying the proposed techniques to regression problems, with accuracy measured by both mean squared error (MSE) and maximum error. While most neural network-based regression methods aim to minimize MSE, often by using it directly as the loss function [27], this approach is suitable only when reducing the average error across the dataset is sufficient. However, it falls short in applications where a single large error could lead to catastrophic failure (e.g., autonomous driving systems or medical diagnostics—see [1], [3]). Our proposed method integrates staged and normalized residual learning with architecture-adaptive modules and explicit $L^{\infty}$ norm error mitigation strategies, enabling highly precise approximations critical in safety-sensitive applications.
In Section 2, we further develop the theory behind gradient boosted neural networks in a more structured way and highlight the resulting approximation error. We write out and test the algorithm on numerous regression examples, as well as investigate the impacts of the number of parameters used. We also compare our proposed method against KANs to understand the trade-offs between developing new methods that take advantage of existing architectures (our proposed method) versus developing completely new architectures (KANs) when trying to improve precision in machine learning. In Section 4 we examine how choice of loss function affects the performance of the technique on both RMSE and $L^{\infty}$ norm error and evaluate it on our chosen examples. In Sections 5 and 6, we introduce the idea of weighted sampling and local patching and evaluate their performance. Finally, in Section 7, we present a boundary-aware training technique accomplished through domain expansion and then provide concluding remarks in Section 8. Overall, we validate the proposed method provides high precision results on a family of benchmark problems. The results demonstrate that the HiPreNet framework effectively improves accuracy measured by both MSE and maximum error. The code used is available at https://github.com/etmulle/residual-training.
@@ 2 High-Precision Neural Networks
Figure 1: Illustration of the HiPreNet training process. Each component is trained sequentially and are combined to form the final model.
@@ 2.1 Algorithmic Formulation of HiPreNet Training
Given a dataset $\{(x_{k},f(x_{k}))\}^{N}_{k=1}$, we will design a neural network based model with high approximation accuracy through an progressive training process based on the residuals. Each neural network will be trained using loss function $L$, which, for regression problems, is typically chosen to be MSE loss:
@@
Thus, the accuracy of the new approximation model $f^{NN}_{i+1}(x)$ keeps increasing with the number of iterations, provided the first neural network approximates the original function with error $e_{0}<1$ and that each residual neural network approximates the normalized residuals with error $\epsilon_{i}<1$. Supposing we have $m$ total refinement networks, this gives a final approximation $f(x)$ as
@@
Figure 2: Illustration of making inferences with the trained model. The same input $x$ is the input to all of the trained networks, where their outputs are summed together to obtain the final prediction $f_{m}^{NN}(x)$.
In practice, several factors can cause the condition of the approximation error of each network to be smaller than one to not be met. Firstly, the residuals may exhibit high-frequency, non-smooth behavior that is hard to approximate with standard architectures, meaning insufficient model capacity in deeper stages may limit residual learning. Additionally, as residuals become increasingly small and oscillatory, numerical precision and optimizer stability may also limit practical convergence. Furthermore, noisy training data or very limited sampling may degrade the estimated residual accuracy, which we do not investigate in this work. In all these cases, the reduction of $L^{\infty}$ norm error may stall or reverse. While we do observe convergence empirically in all of our experiments, formal guarantees remain an open area for future work. The pseudocode of the algorithm is given in Algorithm 1.
@@
Dataset $(\textbf{x},\textbf{f})=\{(x_{k},f(x_{k}))\}^{N}_{k=1}$, trained neural network $f_{0}^{NN}$, tolerance $tol$
@@
$e_{0}\leftarrow\max{|r_{0}(\textbf{x})|}$triangleright estimate of true $e_{0}$
@@
  $f^{NN}_{i}(\textbf{x})\leftarrow f_{i-1}^{NN}+e_{i}\hat{r}_{i}^{NN}(\textbf{x})$
@@
  $e_{i}\leftarrow\max{|r_{i}(\textbf{x})|}$triangleright estimate of true $e_{i}$
@@ 2.2 Preliminary Experiments on a Benchmark Function
Throughout the following sections, we test our algorithm on five different functions from the Feynman dataset first introduced in [30]. Each function is numbered and corresponds to a different dimensionless multi-variable function. These functions correspond to well known physical laws and relations. The functions are written out in Table 1 with their number in the dataset and both their original and dimensionless forms. For this paper, we attempt to learn the dimensionless forms of the functions. The dataset contains 1,000,000 samples for each function, which were all used for training the networks. To test the networks, we randomly generate an additional 1,000,000 validation samples for each function with each variable in the domains given in [30].
@@ I.26.2 $f(n,\theta_{2})=\arcsin{(n\sin\theta_{2})}$ $f(n,\theta_{2})=\arcsin{(n\sin\theta_{2})}$ 2
Table 1: The selected functions from the Feynman dataset with their dimensionless forms.
@@
Each refinement network is optimized with SciPy’s [32] Broyden–Fletcher–Goldfarb–Shanno (BFGS) algorithm [21] for a maximum of 25000 iterations, using a gradient tolerance of 1e-12. The wide and deep fully connected networks are optimized with BFGS for a maximum of 100000 iterations. We use BFGS for training each network due to its strong local convergence properties and robustness in high precision settings. As a quasi-Newton method, BFGS leverages curvature information to guide updates more effectively than first-order methods, enabling more reliable convergence to low error solutions. While computationally more intensive due to requiring more memory and compute time, its performance is well-suited to the relatively small networks used in our framework, making it an ideal choice for achieving the extremely low errors targeted in this work.
@@
Table 2: RMSE for Function I.13.12 using a fully connected network with 5 hidden layers with 11 neurons each (wider), 19 hidden layers with 5 neurons each (deeper), and the described HiPreNet setup with four fully connected networks which each have 5 hidden layers of 5 neurons each,
@@ 2.3 Comparison with Kolmogorov–Arnold Networks (KANs)
For each dataset, an initial predictor neural network as well as three refinement networks were initialized and trained. All networks contain five hidden layers with five neurons each and use tanh as their activation function. Each network was trained using BFGS as the optimizer with 25000 maximum iterations. We run each simulation for three different random seeds and report the results for the seed that provided the lowest validation error.
The initial RMSE and final RMSE after the training the HiPreNets are compared to each other for the different validation datasets in Table 3. We also compare the number of parameters needed for each model. As previously mentioned, in training a HiPreNet, multiple neural networks are trained sequentially and are then combined to create a final model for inference. In Table 3, the “Parameters per Training” column indicates the number of parameters in each individual neural network, while the “Total Model Size” column indicates the total number of parameters in the final model. It is not precisely clear the exact number of parameters used in each KAN from [19], so we report an estimate based on the authors’ presented figures in the “KAN Model Size” column in Table 3.
@@
Function
Initial
RMSE
Final
RMSE
KAN
RMSE
Parameters
per Training
Total
Model Size
KAN
Model Size
@@
Table 3: Comparison of RMSE achieved by HiPreNets and Kolmogorov–Arnold Networks (KANs) across selected functions. Parameter count per each model training and total model sizes are also listed.
Figure 3: (a) RMSE progression across HiPreNet stages. (b) $L^{\infty}$ norm error progression across HiPreNet stages. Model 1 is the initial network and each subsequent model learns to approximate the residuals of the previous model.
@@ 2.4 Empirical Insights from Early Experiments
Table 4: Initial and final $L^{\infty}$ norm errors for each function, showing the change resulting from training a HiPreNet.
@@
Figure 4: Visual comparison of the true function for Function I.6.2, the final model approximation, and the remaining residuals after training.
Figure 5: Surface plots of the residuals from each stage of the HiPreNet training process for Function I.6.2. Note that the axis and colorbar scales between neighboring plots may differ to accurately illustrate the loss landscape from each approach.
@@ 3 Network Width Adaptation in Residual Stages
To confirm that this idea improves the HiPreNet framework, we fix the initial neural network and three residual neural networks to all have a depth of five hidden layers. We then vary the width of the layers for each network to be [5-5-5-5], [20-20-20-20], and [5-10-15-20]. For example, for the [5-10-15-20] case, the initial network will have 5 neurons in each layer, the first refinement network will have 10 neurons in each layer, the second refinement network will have 15, and the third refinement network will have 20. This allows for each additional network to have more capacity to fit the increasingly complex residual structures.
@@
Function
Network
Structure
Final
RMSE
Final $L^{\infty}$
Norm Error
Parameters
per Training
Model
Size
@@
Table 5: Summary of RMSE, $L^{\infty}$ norm error, parameters used for training each network, and total model size across different neuron configurations used in the refinement networks for each function.
Figure 6: Impact of varying neuron counts in successive refinement networks on model validation RMSE and $L^{\infty}$ norm error for all functions.
Figure 7: Residuals at each training stage for Function I.6.2 using a [5-5-5-5] and [5-10-15-20] network configuration. Note that the axis and colorbar scales between neighboring plots may differ to accurately illustrate the loss landscape from each approach.
@@
4 Loss Function Design for $L^{\infty}$ Norm Error Reduction

### v3

@@ Abstract
Deep neural networks are powerful tools for solving nonlinear problems in science and engineering, but training highly accurate models becomes challenging as problem complexity increases. Non-convex optimization and sensitivity to hyperparameters make consistent performance improvement difficult, and traditional approaches prioritize minimizing mean squared error while overlooking the $L^{\infty}$ norm error that is critical in safety-sensitive applications. To address these challenges, we present HiPreNets, a progressive framework for training high-precision neural networks through sequential residual refinements. Starting from an initial network, each stage trains a refinement network on the normalized residuals of the ensemble so far, systematically reducing both average and worst-case error. A key theme throughout the framework is concentrating training effort on high-error regions of the input domain, which we pursue through complementary techniques including loss function design, adaptive data sampling, localized patching, and boundary-aware training. We validate the framework on benchmark regression problems from the Feynman dataset, where it consistently outperforms standard fully connected networks and reported Kolmogorov-Arnold Networks results, with accuracy approaching machine precision depending on select problems. We further apply the framework to learning the flow map of a 20-dimensional power system ODE, which appears to be the highest dimensional problem studied using this class of multistage methods, achieving substantial reductions in both RMSE and $L^{\infty}$ norm error while enabling a surrogate that predicts system state $238\times$ faster than direct numerical simulation.
@@ 1 Introduction
Fully-connected deep neural networks [31, 32, 19] have found widespread use in applications such as computer vision [23] and natural language processing [36, 29], yet they remain underutilized in scientific and engineering domains due to concerns about accuracy and safety [18]. Guaranteeing high accuracy on these problems is difficult because they are often high-dimensional and nonlinear. Much of the existing literature focuses on attaining a “good enough” average error rather than devoting significant effort to drive it closer to machine precision. Additionally, the final absolute maximum error (which we will refer to throughout this work as the $L^{\infty}$ norm error) of a neural network approximation is generally ignored despite worst-case behavior being critical to analyze for sensitive applications [2].
@@
In recent work, [5, 11, 25, 20] have explored the forms of multi-stage or boosted neural networks. Generally in these methods, fully connected neural networks are used as the “weak” models to be combined. These models are less interpretable and more costly to optimize than decision trees, but they are much better at capturing nonlinear patterns, can produce robust outputs, and can more easily be regularized to avoid overfitting. In [38], this approach was taken one step further by recognizing neural networks’ tendency to learn the low-frequency components of functions first, known as the spectral bias of neural networks [30]. As such, the low and high frequency parts of a dataset and the models’ residuals can be analyzed and incorporated into the boosting algorithm to further guide the training process. This analysis is insightful and was shown to work for low dimensional problems, but the spectral analysis it relies upon becomes substantially harder to apply as dimensionality grows. In summary, while these earlier works have applied boosting concepts to neural networks, they typically focused on shallow models or spectral insights that are hard to generalize in practice, none of them analyze methods for reducing $L^{\infty}$ norm error, and none of them have investigated high dimensional systems.
@@
To address these issues, we introduce a progressive training framework based on boosting in which the neural network is composed of multiple components trained sequentially. Each component has significantly lower complexity than the full network. This modularity of the framework also offers great flexibility by allowing for each component network to be trained with a different architecture and loss function that may better fit the residuals it encounters. Crucially, a key theme running throughout the framework is the principle of concentrating training effort on high error regions of the input domain. As we will show, this principle can be followed in multiple complementary ways, such as through loss function design, adaptive data sampling, localized patching, and boundary-aware training, with each targeting the same underlying problem of difficult, localized errors that standard training leaves behind. Together, these techniques allow the training to be guided toward reducing either the MSE or the $L^{\infty}$ norm error, depending on the application’s requirements.
@@
We focus exclusively on applying the proposed techniques to regression problems, with accuracy measured by both mean squared error (MSE) and $L^{\infty}$ norm error. While most neural network-based regression methods aim to minimize MSE, often by using it directly as the loss function [32], this approach is suitable only when reducing the average error across the dataset is sufficient. However, it falls short in applications where a single large error could lead to catastrophic failure (e.g., autonomous driving systems or medical diagnostics—see [1], [3]). Our proposed method integrates staged residual refinement with architecture-adaptive modules and explicit $L^{\infty}$ norm error mitigation strategies, enabling highly precise approximations critical in safety-sensitive applications.
In Section 2, we develop the theory behind gradient boosted neural networks in a structured way and analyze the resulting approximation error. We validate the algorithm on benchmark regression problems, investigate the impact of network width on performance, and compare our method against KANs. In Section 4.1, we examine how loss function choice affects both RMSE and $L^{\infty}$ norm error, introducing a weighted mean squared error loss in line with the mentioned high error concentration principle. In Sections 4.2 and 4.3, we introduce weighted sampling and localized patching as further ways to pursue this principle, each concentrating training resources on regions where the current model performs poorly. In Section 5, we present domain expansion as a boundary-aware technique that addresses a related but distinct source of large errors. In Section 6, we apply the framework to a 20-dimensional power system, where targeted data augmentation provides another way to concentrate on large errors, exploiting simulator access to generate new samples in the high error transient regime. Concluding remarks are given in Section 7.
Overall, we validate the proposed method on a family of benchmark regression problems as well as on learning the flow map of a 20-dimensional power system ODE, which appears to be the first application of this style of multistage neural network framework to a problem that high in dimension. (The highest dimension of a problem studied in a previous work appears to be a 6-dimensional synthetic function regression problem in [25].) Using neural networks to learn flow maps has been an area of great interest in recent years, as summarized in [12], and doing so in moderately high dimensions is challenging due to nonlinearities and the curse of dimensionality, as analyzed in [8]. Our results illustrate that the framework scales to such settings, achieving significant reductions in both RMSE and $L^{\infty}$ norm error relative to a base neural network model, while also enabling a neural network surrogate that predicts the system state 5 seconds forward in time $238\times$ faster than a Forward Euler simulator. The code used is available at https://github.com/etmulle/residual-training.
@@ 2 High-Precision Neural Networks
Figure 1: Illustration of the HiPreNet training process. Starting from an initial network $f_{0}^{NN}$, each subsequent refinement network is trained on the normalized residuals of the ensemble so far, with independent loss function and width. The outputs of all component networks are denormalized and summed to form the final model.
@@ 2.1 Algorithmic Formulation of HiPreNet Training
Given a dataset $\{(x_{k},f(x_{k}))\}^{N}_{k=1}$, we will design a neural network based model with high approximation accuracy through a progressive training process based on the residuals. Each neural network will be trained using loss function $L$, which, for regression problems, is typically chosen to be MSE loss:
@@
Thus, the accuracy of the new approximation model $f^{NN}_{i+1}(x)$ keeps increasing with the number of iterations, provided the first neural network approximates the original function with error $e_{0}<1$ and that each refinement neural network approximates the normalized residuals with error $\epsilon_{i}<1$. Supposing we have $m$ total refinement networks, this gives a final approximation $f(x)$ as
@@
Figure 2: Illustration of making inferences with the trained model. The same input $x$ is the input to all of the trained networks, where their outputs are denormalized and summed together to obtain the final prediction $f_{m}^{NN}(x)$.
As mentioned, the condition $\epsilon_{i}<1$ is necessary for each refinement stage to reduce the approximation error. By the Universal Approximation Theorem (UAT) [19], a sufficiently wide network can approximate any continuous function to arbitrary precision, which guarantees that a network achieving $\epsilon_{i}<1$ exists in principle for each normalized residual $\hat{r}_{i}$. However, three practical factors may prevent the training procedure from finding it. First, the residuals become increasingly oscillatory at later stages, and a network of fixed finite width may lack the capacity to approximate them to the required precision, while the UAT requires unbounded width to guarantee arbitrary accuracy. Second, even if a sufficiently expressive network exists, optimization on a non-convex loss landscape may converge to a local minimum with $\epsilon_{i}\geq 1$. Third, the condition is verified only on the finite training dataset, while the true $L^{\infty}$ norm over the domain may exceed the estimated value $\max|r_{i}|$, particularly in low-density regions. In all these cases, the reduction in $L^{\infty}$ error may stall or reverse. We observe convergence empirically across all experiments presented in this paper, and treat the framework as an empirically validated algorithm with theoretical motivation rather than a method with provable convergence guarantees. Establishing formal guarantees remains an open direction for future work.
The pseudocode of the algorithm is given in Algorithm 1.
@@
Dataset $(\textbf{x},\textbf{f})=\{(x_{k},f(x_{k}))\}^{N}_{k=1}$, trained neural network $f_{0}^{NN}$, tolerance toltol
@@
$e_{0}\leftarrow\max{|r_{0}(\textbf{x})|}$ $\triangleright$ estimate of true $e_{0}$
@@
  $f^{NN}_{i}(\textbf{x})\leftarrow f_{i-1}^{NN}+e_{i-1}\hat{r}_{i-1}^{NN}(\textbf{x})$
@@
  $e_{i}\leftarrow\max{|r_{i}(\textbf{x})|}$ $\triangleright$ estimate of true $e_{i}$
@@ 2.2 Preliminary Experiments on a Benchmark Function
Throughout the following sections, we test our algorithm on five different functions from the Feynman dataset first introduced in [35]. Each function is numbered and corresponds to a different dimensionless multi-variable function. These functions correspond to well known physical laws and relations. The functions are written out in Table 1 with their number in the dataset and both their original and dimensionless forms. For this paper, we attempt to learn the dimensionless forms of the functions. The dataset contains 1,000,000 samples for each function, which were all standardized and used for training the networks. To test the networks, we randomly generate an additional 1,000,000 validation samples for each function with each variable in the domains given in [35].
@@ I.26.2 $f(n,\theta_{2})=\arcsin{(n\sin\theta_{2})}$ $f(n,\theta_{2})=\arcsin{(n\sin\theta_{2})}$ 2
Table 1: Functions selected from the Feynman dataset, shown with their original physical forms, dimensionless forms, and input dimensionality used in experiments.
@@
Each refinement network is optimized with SciPy’s [37] Broyden–Fletcher–Goldfarb–Shanno (BFGS) algorithm [26] for a maximum of 25000 iterations, using a gradient tolerance of 1e-12. The wide and deep fully connected networks are optimized with BFGS for a maximum of 100000 iterations. We use BFGS for training each network due to its strong local convergence properties and robustness in high precision settings. As a quasi-Newton method, BFGS leverages curvature information to guide updates more effectively than first-order methods, enabling more reliable convergence to low error solutions. While computationally more intensive due to scaling as $O(n^{2})$ in memory, its performance is well-suited to the relatively small networks used in our framework, making it an ideal choice for achieving the extremely low errors targeted for these low dimensional functions.
@@
Table 2: Comparison of validation RMSE and parameter counts for wider, deeper, and HiPreNet configurations on Function I.13.12.
@@ 2.3 Comparison with Kolmogorov–Arnold Networks (KANs)
For each Feynman function dataset, an initial predictor neural network as well as three refinement networks were initialized and trained. All networks contain five hidden layers with five neurons each and use tanh as their activation function. Each network was trained using BFGS as the optimizer with 25000 maximum iterations. We run each simulation for three different random seeds and report the results for the seed that provided the lowest validation error.
The initial RMSE and final RMSE after the training the HiPreNets are compared to each other for the different validation datasets in Table 3. We also compare the number of parameters needed for each model. As previously mentioned, in training a HiPreNet, multiple neural networks are trained sequentially and are then combined to create a final model for inference. In Table 3, the “Parameters per Training” column indicates the number of parameters in each individual neural network, while the “Total Model Size” column indicates the total number of parameters in the final model. It is not precisely clear the exact number of parameters used in each KAN from [24], so we report an estimate based on the authors’ presented figures in the “KAN Model Size” column. Following from this, we note that this comparison is approximate, as the KAN results are taken directly from [24] rather than a re-implementation, and differences in aspects such as the training budget and optimizer settings may exist. Nonetheless, the comparison is still informative as an order-of-magnitude benchmark against a recent advanced architecture evaluated on the same dataset.
@@
Function Initial RMSE Final RMSE KAN RMSE Parameters per Iteration Total Model Size KAN Model Size
@@
Table 3: Validation RMSE, per-iteration parameter count, and total model size for HiPreNets versus KANs across all five Feynman functions. KAN results are reproduced from [24]. HiPreNets achieve lower RMSE than the reported KAN results in all cases.
Figure 3: RMSE (a) and $L^{\infty}$ norm error (b) at each stage of HiPreNet training for the five Feynman benchmark functions. Iteration 1 corresponds to the initial network, and each subsequent iteration adds one residual refinement network. While RMSE decreases consistently across all functions, $L^{\infty}$ norm error shows less uniform improvement.
@@ 2.4 Empirical Insights from Early Experiments
Table 4: Initial and final $L^{\infty}$ norm errors for each function, showing the reduction achieved by the HiPreNet training procedure.
@@
Figure 4: Visual comparison of the true function for Function I.6.2, the final model approximation, and the remaining residuals after training. The residuals exhibit peaks concentrated along the domain boundary.
Figure 5: Surface plots of the true and predicted residuals at each stage of HiPreNet training for Function I.6.2, with rows corresponding to successive refinement stages. Note that axis and colorbar scales differ across plots. The overall residual magnitude decreases with each stage, though later residuals exhibit increasingly oscillatory structure that becomes harder to approximate.
@@ 3 Network Width Adaptation in Residual Stages
To confirm that this idea improves the HiPreNet framework, we fix the initial neural network and three refinement neural networks to all have a depth of five hidden layers. We then vary the width of the layers for each network to be [5-5-5-5], [20-20-20-20], and [5-10-15-20]. For example, for the [5-10-15-20] case, the initial network will have 5 neurons in each layer, the first refinement network will have 10 neurons in each layer, the second refinement network will have 15, and the third refinement network will have 20. This allows for each additional network to have more capacity to fit the increasingly complex residual structures.
@@
Function Network Structure Final RMSE Final $L^{\infty}$ Norm Error Parameters per Iteration Model Size
@@
Table 5: Validation RMSE and $L^{\infty}$ norm error for three HiPreNet width schedules ([5-5-5-5], [5-10-15-20], [20-20-20-20]) across all five functions, with per-iteration and total parameter counts (model size).
Figure 6: Validation RMSE (a) and $L^{\infty}$ norm error (b) for all five Feynman functions under three HiPreNet width configurations: uniform [5-5-5-5], linearly increasing [5-10-15-20], and uniform [20-20-20-20]. Linearly increasing widths generally achieve comparable or better accuracy than the [20-20-20-20] configuration at roughly half the total parameter count.
Figure 7: Side-by-side comparison of residuals at each training stage for Function I.6.2, with the [5-5-5-5] configuration (left column) and the [5-10-15-20] configuration (right column). Note that axis and colorbar scales differ across plots. The linearly increasing width configuration achieves substantially smaller residuals by the later stages, reflecting the increased capacity available as residual complexity grows.
@@
4 $L^{\infty}$ Norm Error Reduction
4.1 Loss Function Design for $L^{\infty}$ Norm Error Reduction